1. Under favorable conditions, the human eye can detect light waves with intensities as low as 2.50 × 10−12 W/m2.  At this intensity, what is the average power incident on a pupil of diameter 8.00 mm?

2.  A standing wave has wavenumber 203 rad/m. What is the distance between two adjacent nodes?
Answer #1

I= 101 W/m², diameter, d=8mm Power, Pa I.A = 1012*(4x103)? P=5.024x107 w1 Wave number = 1 I wavelength distance between two